SolveIt is under development
SolveItNCERT · CBSE · NEET

NCERT Exemplar · Class 12 Mathematics Probability

108 questions · 108 still being checked

EXERCISE 13.3 91–100 (part 10 of 11)

  1. Choose the correct answer from the given four options in each of the exercises from $\displaystyle 56$ to 82.

    Exercise 91

    In a college, 30\displaystyle 30% students fail in physics, 25\displaystyle 25% fail in mathematics and 10\displaystyle 10% fail in both. One student is chosen at random. The probability that she fails in physics if she has failed in mathematics is
    (A)
    110\displaystyle \frac{1}{10}
    (B)
    25\displaystyle \frac{2}{5}
    (C)
    920\displaystyle \frac{9}{20}
    (D)
    13\displaystyle \frac{1}{3}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    B
    (B) \(\displaystyle \tfrac25\)\(\displaystyle A\): fails physics, \(\displaystyle B\): fails mathematics. \[P(A)=0.30,\quad P(B)=0.25,\quad P(A\cap B)=0.10 \] \[P(A\mid B)=\frac{P(A\cap B)}{P(B)}=\frac{0.10}{0.25}=\frac{2}{5} \]
  2. Exercise 92

    A and B are two students. Their chances of solving a problem correctly are 13\displaystyle \frac{1}{3} and 14\displaystyle \frac{1}{4}, respectively. If the probability of their making a common error is, 120\displaystyle \frac{1}{20} and they obtain the same answer, then the probability of their answer to be correct is
    (A)
    112\displaystyle \frac{1}{12}
    (B)
    140\displaystyle \frac{1}{40}
    (C)
    13120\displaystyle \frac{13}{120}
    (D)
    1013\displaystyle \frac{10}{13}

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle \tfrac{10}{13}\)Common error is taken as \(\displaystyle \tfrac1{20}\) of the cases in which both are wrong. \[P(\text{both correct})=\tfrac13\cdot\tfrac14=\tfrac1{12} \] \[P(\text{both wrong})=\tfrac23\cdot\tfrac34=\tfrac12 \] \[P(\text{same wrong answer})=\tfrac12\cdot\tfrac1{20}=\tfrac1{40} \] \[P(\text{same answer})=\tfrac1{12}+\tfrac1{40}=\tfrac{13}{120} \] \[P(\text{correct}\mid\text{same answer})=\frac{1/12}{13/120}=\frac{10}{13} \]
  3. Exercise 93

    A box has 100\displaystyle 100 pens of which 10\displaystyle 10 are defective. What is the probability that out of a sample of 5\displaystyle 5 pens drawn one by one with replacement at most one is defective?
    (A)
    (910)5\displaystyle \left(\frac{9}{10}\right)^5
    (B)
    12(910)4\displaystyle \frac{1}{2}\left(\frac{9}{10}\right)^4
    (C)
    12(910)5\displaystyle \frac{1}{2}\left(\frac{9}{10}\right)^5
    (D)
    (910)5+12(910)4\displaystyle \left(\frac{9}{10}\right)^5+\frac{1}{2}\left(\frac{9}{10}\right)^4

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    D
    (D) \(\displaystyle \left(\tfrac{9}{10}\right)^5+\tfrac12\left(\tfrac{9}{10}\right)^4\)Drawing with replacement gives a binomial distribution: \[n=5,\quad p=\tfrac{10}{100}=\tfrac1{10},\quad q=\tfrac{9}{10} \] \[P(X\le1)=P(X=0)+P(X=1)=\binom50q^5+\binom51pq^4 \] \[=\left(\tfrac{9}{10}\right)^5+5\cdot\tfrac1{10}\left(\tfrac{9}{10}\right)^4=\left(\tfrac{9}{10}\right)^5+\tfrac12\left(\tfrac{9}{10}\right)^4 \]
  4. State True or False for the statements in each of the Exercises $\displaystyle 94$ to 103.

    Exercise 94

    Let P(A)>0\displaystyle \mathrm{P}(\mathrm{A})>0 and P(B)>0\displaystyle \mathrm{P}(\mathrm{B})>0. Then A and B can be both mutually exclusive and independent.

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    False
    False. The two conditions give contradictory values of \(\displaystyle P(A\cap B)\). \[A,B\ \text{mutually exclusive}:\quad P(A\cap B)=0 \] \[A,B\ \text{independent}:\quad P(A\cap B)=P(A)\,P(B)>0 \quad\text{since } P(A)>0,\ P(B)>0 \]
  5. Exercise 95

    If A and B are independent events, then A′\displaystyle \mathrm{A}^{\prime} and B′\displaystyle \mathrm{B}^{\prime} are also independent.

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    True
    True. \[P(A\cap B)=P(A)\,P(B) \quad\text{(given)} \] \[P(A'\cap B')=P\big((A\cup B)'\big)=1-P(A\cup B) \quad\text{(De Morgan)} \] \[=1-P(A)-P(B)+P(A\cap B) \] \[=1-P(A)-P(B)+P(A)\,P(B) \] \[=\big(1-P(A)\big)\big(1-P(B)\big)=P(A')\,P(B') \]
  6. Exercise 96

    If A and B are mutually exclusive events, then they will be independent also.

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    False
    False. Mutually exclusive events have \(\displaystyle P(A\cap B)=0\), which differs from \(\displaystyle P(A)P(B)\) unless one of them is impossible.Toss a coin: \(\displaystyle A\) = head, \(\displaystyle B\) = tail. \[P(A\cap B)=0\ne\tfrac14=P(A)\,P(B) \]
  7. Exercise 97

    Two independent events are always mutually exclusive.

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    False
    False. Independent events of positive probability cannot be disjoint: \[P(A\cap B)=P(A)\,P(B)>0 \] \[\text{mutually exclusive} \Rightarrow P(A\cap B)=0 \] \[P(A)=P(B)=\tfrac12 \Rightarrow P(A\cap B)=\tfrac14\neq 0 \]
  8. Exercise 98

    If A and B are two independent events then P(A and B)=P(A).P(B)\displaystyle \mathrm{P}(\mathrm{A} \text { and } \mathrm{B})=\mathrm{P}(\mathrm{A}) . \mathrm{P}(\mathrm{B}).

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    True
    True. This is the multiplication rule that defines independence: \[P(A\cap B)=P(A)\,P(B) \]
  9. Exercise 99

    Another name for the mean of a probability distribution is expected value.

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    True
    True. The mean of a probability distribution is its expected value: \[E(X)=\sum_{i} x_i\,p_i \]
  10. Exercise 100

    If A and B′\displaystyle \mathrm{B}^{\prime} are independent events, then P(A′∪B)=1−P(A)P(B′)\displaystyle \mathrm{P}\left(\mathrm{A}^{\prime} \cup \mathrm{B}\right)=1-\mathrm{P}(\mathrm{A}) \mathrm{P}\left(\mathrm{B}^{\prime}\right)

    Matches the book, not yet reviewed

    This working reaches the answer NCERT prints. It has not yet been read through by hand.

    NCERT’s answer
    True
    True. \(\displaystyle A'\cup B\) is the complement of \(\displaystyle A\cap B'\) (De Morgan), and \(\displaystyle A,\,B'\) are independent: \[P(A'\cup B)=1-P(A\cap B') \] \[P(A'\cup B)=1-P(A)\,P(B') \]